OpenStudy (anonymous):

No idea how to do this.... Find all the solutions of the trigonometric equation on the interval 0 < x < 2pi

OpenStudy (anonymous):

\[2 \sin^2 x - \sin x -1 =0\]

OpenStudy (primeralph):